OpenStudy (anonymous):

Solve the following formula for x y = 5 mx -6 b x=

OpenStudy (littlebird):

Are the m and b supposed to be there?

OpenStudy (anonymous):

yes.

OpenStudy (abbles):

x = (y + 6)/5 is your answer

OpenStudy (anonymous):

Solve the following formula for x y = 6 mx -3 b x=

OpenStudy (abbles):

You simply add 6 to both sides and then divide by 5.

OpenStudy (anonymous):

and this one ?

OpenStudy (anonymous):

@Abbles

OpenStudy (littlebird):

If I were to solve that equations as is.....

OpenStudy (abbles):

x = (y + 3)/6

OpenStudy (littlebird):

I still think the m and b are weird

OpenStudy (abbles):

@Littlebird I think the m and b are just placeholders, because it is in y = mx + b form

OpenStudy (anonymous):

it says its wrong that they both need paranthesis

OpenStudy (abbles):

Are you saying that the parenthesis aren't needed or that the answer was wrong? I put the parenthesis so that you would know what was being divided by what, but on paper of course (or on an online course) the parenthesis might not be necessary.

OpenStudy (the_fizicx99):

Are you sure it's y = 5mx - 6b? in y = mx + b they're place holders... If it was on purpose then subtract 6b from both sides, and divide both sides by 5m \(\ \dfrac{y + 6b}{5m} = x \)